OpenStudy (anonymous):

In walking across a carpet, you acquire a net negative charge of 50 µC. How many electrons do you have?

OpenStudy (anonymous):

Hi you can use this: \[q=n e\] where q=50*10^-6 (C) & e=1.6*10^-19 (C)

OpenStudy (anonymous):

Thanks